# Einstein metrics on homogeneous superspaces

### v1

@@ 2.3 Harish-Chandra pairs
Given a Lie supergroup $G$, we can define a Harish-Chandra pair $(G_{0},\mathfrak{g})$ by taking $G_{0}$ to be the underlying reduced Lie group of $G$ and $\mathfrak{g}$ to be the Lie superalgebra of $G$. It is well known that the category of (real) Harish-Chandra pairs is equivalent to the category of (real) Lie supergroups; see, e.g., [CCF11, Theorem 7.4.5]. Thus, a Lie supergroup only depends on the underlying ordinary Lie group and Lie superalgebra, together with the compatibility conditions given in Definition 2.1.
@@
$\operatorname{SL}(m|n)_{0}=\{(A,B)\in\operatorname{GL}(m)\times\operatorname{GL}(n)\mid\mathrm{det}(A)=\mathrm{det}(B)>0\}$
@@
$\mathfrak{sl}(m|n)=\{X\in\mathfrak{gl}(m|n)\mid\operatorname{Str}(X)=0\}.$
A representation of a Harish-Chandra pair $G=(G_{0},\mathfrak{g})$ on a vector superspace $V$ consists of a representation of $G_{0}$ on $V$ and one of $\mathfrak{g}$ on $V$, such that the differential of the representation of $G_{0}$ agrees with the representation of the even part $\mathfrak{g}_{\bar{0}}$ on $V$.
@@ 2.4 Homogeneous superspaces
The super-structure sheaf $\mathcal{O}_{G/K}$ on $G_{0}/K_{0}$ is defined as follows [Kos77, Theorem 3.9]. Consider the projection $\mathrm{pr}_{1}:G\times K\to G$ onto the first factor, and the morphism $\psi:G\times K\to G$ induced by the multiplication $m$. For any open subset $U\subseteq G_{0}/K_{0}$, define the set of $K$-invariant sections by
$\mathcal{O}_{G/K}(U):=\{f\in\mathcal{O}_{G}(\pi_{0}^{-1}(U))\mid\psi^{*}f=\mathrm{pr}_{1}^{*}f\}.$
@@
referred to as a homogeneous superspace. Let $\pi:G\to G/K$ be the corresponding canonical surjective morphism of supermanifolds, where the sheaf morphism is given by inclusion. The differential of $\pi$ at the identity $e\in G$ yields a surjective linear map $\mathfrak{g}\cong T_{e}G\to T_{K}(G/K)$ with kernel $\mathfrak{k}$, the Lie superalgebra of $K$. This induces a canonical identification: $T_{K}(G/K)\cong\mathfrak{g}/\mathfrak{k}$.
A Harish-Chandra pair $G=(G_{0},\mathfrak{g})$ acts on a supermanifold $M$ if the reduced Lie group $G_{0}$ acts on $|M|$, and there exists a Lie superalgebra anti-morphism from $\mathfrak{g}$ to the Lie superalgebra $\mathcal{T}_{M}(M)$ of vector fields on $M$ such that its restriction to $\mathfrak{g}_{\bar{0}}$ agrees with the differential of the $G_{0}$-action at the identity [DM99, §3.8]. Indeed, with $\alpha:G\times M\to M$ denoting the Lie supergroup action, the anti-morphism is given by
@@
$\mathfrak{g}\to T_{p}M,\qquad X\mapsto\mathrm{ev}_{p}\circ(X_{e}\otimes I)\circ\alpha^{*},$ (2.4)
is surjective [Goe08, Section 4.11]. Here, $\mathrm{ev}_{p}$ is the evaluation of vector fields at $p$.
A supermanifold $M$ is said to be $G$-homogeneous if the Lie supergroup $G$ acts on it transitively. Every $G$-homogeneous supermanifold $M$ is isomorphic to $G/G_{p}$ with $G_{p}$ being the isotropy subsupergroup of $G$ at $p\in|M|$ [Kos77, Propostion 3.10.3]. By definition, the isotropy subsupergroup $G_{p}$ is the Harish-Chandra pair $((G_{0})_{p},\mathfrak{g}_{p})$, where $(G_{0})_{p}$ denotes the ordinary isotropy subgroup of $G_{0}$ at $p$, and $\mathfrak{g}_{p}$ consists of all left-invariant vector fields $X$ on $G$ such that the infinitesimal action of $X$ at $p$ is trivial; see [Goe08, Section 4.11] and [CCF11, §8.4].
@@ 3 Curvature
3.1 Graded Riemannian metrics and connections
Following [Goe08], a graded Riemannian metric on a supermanifold $M=(|M|,\mathcal{O}_{M})$ is a supersymmetric non-degenerate even $\mathcal{O}_{M}$-linear morphism of sheaves $\langle-,-\rangle:\mathcal{T}_{M}\otimes\mathcal{T}_{M}\to\mathcal{O}_{M}$. For each $p\in|M|$, this induces a scalar superproduct on the tangent superspace $T_{p}M$,
$\langle-,-\rangle_{p}:T_{p}M\times T_{p}M\to\mathbb{R},$
that is symmetric on $(T_{p}M)_{\bar{0}}\times(T_{p}M)_{\bar{0}}$, skew-symmetric on $(T_{p}M)_{\bar{1}}\times(T_{p}M)_{\bar{1}}$, and zero on $(T_{p}M)_{\bar{0}}\times(T_{p}M)_{\bar{1}}$ and $(T_{p}M)_{\bar{1}}\times(T_{p}M)_{\bar{0}}$. By the non-degeneracy of the metric, $\dim(M_{\bar{1}})$ must be even. Moreover, the restriction of $\langle-,-\rangle_{p}$ to $(T_{p}M)_{\bar{0}}$ may be indefinite, in which case it gives rise to a pseudo-Riemannian metric on the reduced manifold $M_{0}$. A supermanifold equipped with a graded Riemannian metric is called a Riemannian supermanifold.
A Levi-Civita connection satisfies
$X\,\langle Y,Z\rangle$ $=\langle\nabla_{X}Y,Z\rangle+(-1)^{[X][Y]}\langle Y,\nabla_{X}Z\rangle,$ (3.1)
@@ A Levi-Civita connection satisfies
for all vector fields $X,Y,Z$ on $M$, meaning that the connection is metric and torsion-free. The metric condition (3.1) is trivial if $[X]+[Y]+[Z]=\bar{1}$, noting that $X\,\langle Y,Z\rangle=0$ for all $Y,Z$ if $[X]=\bar{1}$. As in the non-super case, the Levi-Civita connection is unique and determined by a Koszul formula.
@@
[MS96, Theorem 4.2] Let $(M,\langle-,-\rangle)$ be a Riemannian supermanifold. Then, there exists a unique torsion-free metric connection $\nabla$ on $M$, given implicitly by the Koszul formula
$2\,\langle\nabla_{X}Y,Z\rangle$ $=X\,\langle Y,Z\rangle+(-1)^{[X][Y]}Y\,\langle X,Z\rangle-(-1)^{([X]+[Y])[Z]}Z\,\langle X,Y\rangle$
$\quad-\langle X,[Y,Z]\rangle-(-1)^{[X][Y]}\langle Y,[X,Z]\rangle+(-1)^{([X]+[Y])[Z]}\langle Z,[X,Y]\rangle,$
@@
Let $G=(G_{0},\mathfrak{g})$ be a Lie supergroup. A graded Riemannian metric $\langle$-,$\rangle$ - , - ⟩ on $G$ is called left-invariant if $\langle X,Y\rangle$ is a constant function on $G$ for all left-invariant vector fields $X,Y$. We identify the Lie superalgebra $\mathfrak{g}$ of left-invariant vector fields with the tangent space $T_{e}G$ at the identity $e\in G$. For any scalar superproduct $\langle-,-\rangle_{e}$ on $\mathfrak{g}$ and $X,Y\in\mathfrak{g}$, we may then view the real number $\langle X,Y\rangle_{e}$ as a constant function on $G$. For ease of notation, we usually omit the subscript $e$. As in the non-super case, the left-invariant graded Riemannian metrics on $G$ are in natural bijection with the scalar superproducts on $\mathfrak{g}$.
Lemma 3.2.
Let $\langle$-,$\rangle$ - , - ⟩ be a left-invariant graded Riemannian metric on $(G_{0},\mathfrak{g})$. Then,
$2\,\langle\nabla_{X}Y,Z\rangle=\langle[X,Y],Z\rangle-\langle X,[Y,Z]\rangle-(-1)^{[X][Y]}\langle Y,[X,Z]\rangle,\qquad X,Y,Z\in\mathfrak{g}.$
Proof.
Since $\langle$-,$\rangle$ - , - ⟩ is left-invariant, $X\,\langle Y,Z\rangle=0$ for all $X,Y,Z\in\mathfrak{g}$, so the result follows from Proposition 3.1. ∎
A scalar superproduct $\langle$-,$\rangle$ - , - ⟩ on $\mathfrak{g}$ is called $\operatorname{Ad}_{G}$-invariant if it is both $\operatorname{Ad}_{G_{0}}$- and $\operatorname{ad}_{\mathfrak{g}}$-invariant, i.e., for every $X,Y,Z\in\mathfrak{g}$ and every $a\in G_{0}$, we have
Note that if $G$ is connected, then $\operatorname{ad}_{\mathfrak{g}}$-invariance implies $\operatorname{Ad}_{G_{0}}$-invariance. A left-invariant graded Riemannian metric on $G$ is bi-invariant if and only if the corresponding scalar product on $\mathfrak{g}$ is $\operatorname{Ad}_{G}$-invariant [Goe08, Theorem 2].
@@
$X\,\langle Y,Z\rangle=\langle[X,Y],Z\rangle+(-1)^{[X][Y]}\langle Y,[X,Z]\rangle,$ (3.3)
where $Y,Z$ are any vector fields, and we note that the set of graded Killing vector fields forms a Lie superalgebra. A graded Riemannian metric on a $G$-homogeneous superspace $M$ is $G$-invariant if (i) every element of $G_{0}$ acts on $M$ by isometries, and (ii) the image of the Lie superalgebra anti-homomorphism $\mathfrak{g}\to\mathcal{T}_{M}(M)$ is contained in the Lie superalgebra of graded Killing vector fields.
Theorem 3.3.
[Goe08, Theorem 3] Let $G=(G_{0},\mathfrak{g})$ be a Lie supergroup and $K=(K_{0},\mathfrak{k})$ a closed Lie subsupergroup of $G$. The map $\langle-,-\rangle\mapsto\langle-,-\rangle_{K}$ is a bijective correspondence between $G$-invariant graded Riemannian metrics on $G/K$ and $\operatorname{Ad}_{K}$-invariant scalar superproducts on $T_{K}(G/K)$.
Using (3.1) and (3.2), a vector field $X$ is seen to be a graded Killing vector field if and only if
$\langle\nabla_{Y}X,Z\rangle+(-1)^{[X][Z]}\langle Y,\nabla_{Z}X\rangle=0$ (3.4)
@@
Let $X,Y,Z$ be graded Killing vector fields on a Riemannian supermanifold $M$. Then,
$2\,\langle\nabla_{X}Y,Z\rangle=\langle X,[Y,Z]\rangle+X\,\langle Y,Z\rangle=\langle X,[Y,Z]\rangle+\langle[X,Y],Z\rangle+(-1)^{[X][Y]}\langle Y,[X,Z]\rangle.$
3.2 Riemannian, Ricci and scalar curvatures
Given a Riemannian supermanifold $(M,\langle-,-\rangle)$, let $\{X_{1},\dots,X_{d}\}$ be an ordered homogeneous basis for the tangent space $T_{p}(M)$ at $p\in M$, with $\{\widebar{X}_{1},\ldots,\widebar{X}_{d}\}$ the corresponding right dual basis. Thus, if $\eta$ is the matrix representation of $\langle-,-\rangle_{p}$ relative to the ordered basis, then $\widebar{X}_{i}=\sum_{j=1}^{d}(\eta^{-1})_{ji}X_{j}$, and $[\widebar{X}_{i}]=[X_{i}]$ for all $i$. Note that, for any $X\in T_{p}(M)$, we have
@@ 3.2 Riemannian, Ricci and scalar curvatures
In the following, we will be omitting references to the point $p$.
Following [DeW92], the Riemann curvature tensor is defined by
where $X,Y,Z$ are any vector fields on $M$. Using this, the Ricci curvature is defined by
$\operatorname{Ric}(X_{i},X_{k}):=\sum_{j=1}^{d}(-1)^{[X_{j}]+[X_{j}][X_{k}]}\langle R(X_{i},X_{j})X_{k},\widebar{X}_{j}\rangle.$ (3.6)
It follows that $\operatorname{Ric}(-,-)$ is even, so $\operatorname{Ric}(X_{i},X_{j})=0$ if $[X_{i}]\neq[X_{j}]$. Finally, the scalar curvature is defined as the supertrace of the Ricci curvature:
$S:=\sum_{i=1}^{d}(-1)^{[X_{i}]}\operatorname{Ric}(X_{i},\widebar{X}_{i}).$ (3.7)

### v2

@@ 2.3 Harish-Chandra pairs
Given a Lie supergroup $G$, we can define a Harish-Chandra pair $(G_{0},\mathfrak{g})$ by taking $G_{0}$ to be the underlying reduced Lie group of $G$, and $\mathfrak{g}$ to be the Lie superalgebra of $G$. It is well known that the category of (real) Harish-Chandra pairs is equivalent to the category of (real) Lie supergroups; see, e.g., [CCF11, Theorem 7.4.5]. Thus, a Lie supergroup only depends on the underlying ordinary Lie group and Lie superalgebra, together with the compatibility conditions given in Definition 2.1.
@@
$\operatorname{SL}(m|n)_{0}=\{(A,B)\in\operatorname{GL}(m)\times\operatorname{GL}(n)\,|\,\mathrm{det}(A)=\mathrm{det}(B)>0\}$
@@
$\mathfrak{sl}(m|n)=\{X\in\mathfrak{gl}(m|n)\,|\,\operatorname{Str}(X)=0\}.$
A representation of a Harish-Chandra pair $G=(G_{0},\mathfrak{g})$ on a vector superspace $V$ consists of a representation of $G_{0}$ on $V$ and one of $\mathfrak{g}$ on $V$, such that the differential of the representation of $G_{0}$ agrees with the restriction to $\mathfrak{g}_{0}$ of the $\mathfrak{g}$-representation.
@@ 2.4 Homogeneous superspaces
Following [Kos77, Theorem 3.9], the super-structure sheaf $\mathcal{O}_{G/K}$ on $G_{0}/K_{0}$ is defined as follows. Let $\mathrm{pr}_{1}:G\times K\to G$ denote the projection onto the first factor, and $\psi:G\times K\to G$ the morphism induced by the multiplication $m$. For any open subset $U\subseteq G_{0}/K_{0}$, define the set of $K$-invariant sections by
$\mathcal{O}_{G/K}(U):=\{f\in\mathcal{O}_{G}(\pi_{0}^{-1}(U))\,|\,\psi^{*}f=\mathrm{pr}_{1}^{*}f\}.$
@@
referred to as a homogeneous superspace. Let $\pi:G\to G/K$ be the corresponding canonical surjective homomorphism of supermanifolds, where the sheaf morphism is given by inclusion. The differential of $\pi$ at the identity $e\in G$ yields a surjective linear map $\mathfrak{g}\cong T_{e}G\to T_{K}(G/K)$ with kernel $\mathfrak{k}$, the Lie superalgebra of $K$. This induces a canonical identification: $T_{K}(G/K)\cong\mathfrak{g}/\mathfrak{k}$.
A Harish-Chandra pair $G=(G_{0},\mathfrak{g})$ acts on a supermanifold $M$ if the reduced Lie group $G_{0}$ acts on $|M|$, and there exists a Lie superalgebra anti-homomorphism from $\mathfrak{g}$ to the Lie superalgebra $\mathcal{T}_{M}(M)$ of vector fields on $M$ such that its restriction to $\mathfrak{g}_{\bar{0}}$ agrees with the differential of the $G_{0}$-action at the identity [DM99, §3.8]. Indeed, with $\alpha:G\times M\to M$ denoting the Lie supergroup action, the anti-homomorphism is given by
@@
$\mathfrak{g}\to T_{p}M,\qquad X\mapsto\mathrm{ev}_{p}\circ(X_{e}\otimes 1)\circ\alpha^{*},$ (2.6)
is surjective [Goe08, Section 4.11].
A supermanifold $M$ is said to be $G$-homogeneous if the Lie supergroup $G$ acts on it transitively. Every $G$-homogeneous supermanifold $M$ is isomorphic to $G/G_{p}$ with $G_{p}$ being the isotropy subsupergroup of $G$ at $p\in|M|$ [Kos77, Proposition 3.10.3]. By definition, the isotropy subsupergroup $G_{p}$ is the Harish-Chandra pair $((G_{0})_{p},\mathfrak{g}_{p})$, where $(G_{0})_{p}$ denotes the ordinary isotropy subgroup of $G_{0}$ at $p$, and $\mathfrak{g}_{p}$ consists of all left-invariant vector fields $X$ on $G$ such that the infinitesimal action of $X$ at $p$ is trivial; see [Goe08, Section 4.11] and [CCF11, §8.4]. The isotropy representation of $G_{p}$ on $T_{p}M$ is equivalent to the adjoint representation on $\mathfrak{g}/\mathfrak{g}_{p}$ via the natural isomorphism $\mathfrak{g}/\mathfrak{g}_{p}\cong T_{p}M$ [Goe08, Proposition 12].
@@ 3 Curvature
3.1 Graded Riemannian metrics and curvatures
Let $V=V_{\bar{0}}\oplus V_{\bar{1}}$ be a vector superspace over $\mathbb{R}$. A scalar superproduct on $V$ is a non-degenerate, supersymmetric, even bilinear form $\langle-,-\rangle:V\times V\rightarrow\mathbb{R}$. By the non-degeneracy, $\dim(V_{\bar{1}})$ must be even. The supersymmetry implies that $\langle-,$-$\rangle$ is symmetric on $V_{\bar{0}}\times V_{\bar{0}}$, and skew-symmetric on $V_{\bar{1}}\times V_{\bar{1}}$, while the evenness implies that it is zero on $V_{\bar{0}}\times V_{\bar{1}}$ and $V_{\bar{1}}\times V_{\bar{0}}$. Moreover, the restriction of $\langle-,$-$\rangle$ to $V_{\bar{0}}$ may be indefinite.
Definition 3.1.
[Goe08] A graded Riemannian metric on a supermanifold $M=(|M|,\mathcal{O}_{M})$ is an $\mathcal{O}_{M}$-linear morphism of sheaves $g:\mathcal{T}_{M}\otimes\mathcal{T}_{M}\to\mathcal{O}_{M}$ such that for any $X,Y\in\mathcal{T}_{M}(U)$, where $U$ is any open subset of $|M|$,
  1. (1)
    (non-degeneracy) the map $X\mapsto g(X,-)$ is an isomorphism $\mathcal{T}_{M}\rightarrow\mathcal{T}_{M}^{*}$;
  2. (2)
    (supersymmetry) $g(X,Y)=(-1)^{[X][Y]}g(Y,X)$;
  3. (3)
    (evenness) $[g(X,Y)]=[X]+[Y]$.
For each $p\in|M|$, the sheaf morphism $g$ induces a scalar superproduct $\langle-,-\rangle_{p}:=g_{p}$ on the tangent superspace $T_{p}M$,
$\langle-,-\rangle_{p}:T_{p}M\times T_{p}M\to\mathbb{R}.$
If the restriction of $\langle-,-\rangle_{p}$ to $(T_{p}M)_{\bar{0}}$ is indefinite, it induces a pseudo-Riemannian metric on the reduced manifold $M_{0}$. We stress that the scalar superproduct $\langle-,-\rangle_{p}$ does not determine the graded Riemannian metric unless the odd dimension of $M$ is zero. A supermanifold $M$ equipped with a graded Riemannian metric $g$ is called a Riemannian supermanifold, and is denoted by $(M,g)$.
An affine connection on a supermanifold $M$ is an even morphism $\nabla:\mathcal{T}_{M}\otimes_{\mathbb{R}}\mathcal{T}_{M}\rightarrow\mathcal{T}_{M}$ of sheaves of vector superspaces such that, for every open subset $U\subseteq|M|$,
$\nabla_{fX}Y=f\nabla_{X}Y,\qquad\nabla_{X}(fY)=X(f)Y+(-1)^{[X][f]}f\nabla_{X}Z,$
where $f\in\mathcal{O}_{M}(U)$ and $X,Y\in\mathcal{T}_{M}(U)$. If $(M,g)$ is a Riemannian supermanifold, there is a unique affine connection, called the Levi-Civita connection, satisfying
$Xg(Y,Z)$ $=g(\nabla_{X}Y,Z)+(-1)^{[X][Y]}g(Y,\nabla_{X}Z),$ (3.1)
@@ 3.1 Graded Riemannian metrics and curvatures
for all vector fields $X,Y,Z$ on $M$, meaning that the connection is metric and torsion-free. As in the non-super case, the Levi-Civita connection is determined by the following super analogue of the Koszul formula.
@@
[MS96, Theorem 4.2] Let $(M,g)$ be a Riemannian supermanifold. Then, the Levi-Civita connection $\nabla$ is given implicitly by the Koszul formula
$2\,g(\nabla_{X}Y,Z)$ $=Xg(Y,Z)+(-1)^{[X][Y]}Yg(X,Z)-(-1)^{([X]+[Y])[Z]}Zg(X,Y)$
$\quad-g(X,[Y,Z])-(-1)^{[X][Y]}g(Y,[X,Z])+(-1)^{([X]+[Y])[Z]}g(Z,[X,Y]),$
@@
$Xg(Y,Z)=g([X,Y],Z)+(-1)^{[X][Y]}g(Y,[X,Z]),$ (3.3)
where $Y,Z$ are any vector fields, and we note that the set of graded Killing vector fields forms a Lie superalgebra. Using (3.1) and (3.2), a vector field $X$ is seen to be a graded Killing vector field if and only if
$g(\nabla_{Y}X,Z)+(-1)^{[X][Z]}g(Y,\nabla_{Z}X)=0$ (3.4)
@@
Let $X,Y,Z$ be graded Killing vector fields on a Riemannian supermanifold $(M,g)$. Then,
$2g(\nabla_{X}Y,Z)=g(X,[Y,Z])+Xg(Y,Z)=g(X,[Y,Z])+g([X,Y],Z)+(-1)^{[X][Y]}g(Y,[X,Z]).$
Following [DeW92], the Riemann curvature tensor on a Riemannian supermanifold $(M,g)$ is defined by
where $X,Y,Z\in\mathcal{T}_{M}(U)$ are any vector fields and $U$ is any open subset of $|M|$. Furthermore, the Ricci curvature tensor is defined as
$\mathrm{Ric}(Y,Z)=\mathrm{Str}\big(X\mapsto(-1)^{[X][Z]}R(Y,X)Z\big),$
where the supertrace $\mathrm{Str}$ is taken over the $\mathcal{O}_{M}(U)$-linear superspace $\mathcal{T}_{M}(U)$. Similarly, the scalar curvature is defined as the supertrace of the Ricci curvature tensor with respect to the metric $g$.
In the remainder of this paper, we will work with these curvature tensors for homogeneous superspaces. To present the curvature tensors evaluated at a point $p\in|M|$, let $\{X_{1},\dots,X_{d}\}$ be an ordered homogeneous basis for the tangent superspace $T_{p}(M)$, with $\{\widebar{X}_{1},\ldots,\widebar{X}_{d}\}$ the corresponding right dual basis. For a vector $X\in T_{p}(M)$, we thus have
@@
At $p\in|M|$, the Ricci curvature is given by
$\operatorname{Ric}(X_{i},X_{k})_{p}:=\sum_{j=1}^{d}(-1)^{[X_{j}]+[X_{j}][X_{k}]}\langle R(X_{i},X_{j})X_{k},\widebar{X}_{j}\rangle_{p}.$ (3.6)
As $\operatorname{Ric}(-,-)_{p}$ is even, we have $\operatorname{Ric}(X_{i},X_{j})_{p}=0$ if $X_{i}]\neq[X_{j}]$. Finally, the scalar curvature at $p$ is
$S(p):=\sum_{i=1}^{d}(-1)^{[X_{i}]}\operatorname{Ric}(X_{i},\widebar{X}_{i})_{p}.$ (3.7)
3.2 Invariant metrics
Let $G=(G_{0},\mathfrak{g})$ be a Lie supergroup. A graded Riemannian metric $g$ on $G$ is called left-invariant if $g(X,Y)$ is a constant function on $G$ for all left-invariant vector fields $X,Y$. Recall that we identify the Lie superalgebra $\mathfrak{g}$ of left-invariant vector fields with the tangent space $T_{e}G$ at the identity $e\in G$ via the isomorphism $X\mapsto X_{e}$. Similarly to the non-super case, the left-invariant graded Riemannian metrics $g$ on $G$ are in natural bijection with the scalar superproducts $\langle-,-\rangle:=g_{e}$ on $\mathfrak{g}\cong T_{e}G$ [Goe08, Section 4.9].
Lemma 3.4.
Let $g$ be a left-invariant graded Riemannian metric on $(G_{0},\mathfrak{g})$. Then,
$2\,g(\nabla_{X}Y,Z)=g([X,Y],Z)-g(X,[Y,Z])-(-1)^{[X][Y]}g(Y,[X,Z]),\qquad X,Y,Z\in\mathfrak{g}.$
Proof.
Since $g$ is left-invariant, $Xg(Y,Z)=0$ for all $X,Y,Z\in\mathfrak{g}$, so the result follows from Proposition 3.2. ∎
A scalar superproduct $\langle-,$-$\rangle$ on a Lie superalgebra $\mathfrak{g}\cong T_{e}G$ is called $\operatorname{Ad}_{G}$-invariant if it is both $\operatorname{Ad}_{G_{0}}$- and $\operatorname{ad}_{\mathfrak{g}}$-invariant, i.e., for every $X,Y,Z\in\mathfrak{g}$ and every $a\in G_{0}$, we have
Note that if $G$ is connected, then $\operatorname{ad}_{\mathfrak{g}}$-invariance implies $\operatorname{Ad}_{G_{0}}$-invariance.
Definition 3.5.
A graded Riemannian metric on a $G$-homogeneous superspace $M$ is $G$-invariant if (i) every element of $G_{0}$ acts on $M$ by isometries, and (ii) the image of the anti-homomorphism (2.5) is contained in the Lie superalgebra of graded Killing vector fields.
Theorem 3.6.
[Goe08, Theorem 3] Let $G=(G_{0},\mathfrak{g})$ be a Lie supergroup and $K=(K_{0},\mathfrak{k})$ a closed Lie subsupergroup of $G$. The map $g\mapsto g_{K}$ is a bijective correspondence between $G$-invariant graded Riemannian metrics on $G/K$ and $\operatorname{Ad}_{K}$-invariant scalar superproducts on the tangent superspace $T_{K}(G/K)$ at the base point $K$.
Theorem 3.6 implies that every $G$-invariant graded Riemannian metric $g$ on $G/K$ is uniquely determined by its evaluation $g_{K}$ at the base point $K$. Using the method in the proof of this result in [Goe08], one shows that $\mathrm{Ric}(g)$ on $G/K$ is uniquely determined by $\mathrm{Ric}(g)_{K}$, even though the Ricci tensor may be degenerate. Here, $\mathrm{Ric}(g)_{K}$ is a supersymmetric, $\mathrm{Ad}_{K}$-invariant even bilinear form on $T_{K}(G/K)$. More generally [San10, Theorem 4.16], there is a bijective correspondence between $G$-invariant tensor fields of type $(r,s)$ on $G/K$ and isotropy-invariant tensors of type $(r,s)$ on $T_{K}(G/K)$, via evaluation at $K$. In our case, both $g$ and $\mathrm{Ric}(g)$ are $G$-invariant tensor fields of type $(0,2)$. Thus, to solve the Einstein equation $\mathrm{Ric}(g)=cg$ on a homogeneous Riemannian supermanifold $(G/K,g)$, it is necessary and sufficient to solve the corresponding equation at the base point $K$, that is, to find $g_{K}$ satisfying $\mathrm{Ric}(g)_{K}=cg_{K}$. The bijective correspondence ensures that a solution at $K$ uniquely extends to a $G$-invariant solution on the entire supermanifold.